OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x+y=-1 x- 3y=11 A. (2, −3) B. (−2, 1) C. (5, −2) D. (−6, 5)

OpenStudy (anonymous):

will medal and fan

OpenStudy (lukeskylicker):

So. This Pre-Algebra?

OpenStudy (anonymous):

ya im really bad at math

OpenStudy (lukeskylicker):

No your not. It's algebra few people actually understand it.

OpenStudy (lukeskylicker):

So their are two ways to do this. The cheaty way and the way you'll use if this was real life.

OpenStudy (lukeskylicker):

Let's do the later.

OpenStudy (anonymous):

okay

OpenStudy (lukeskylicker):

So do you remember how to solve a system of equations?

OpenStudy (anonymous):

no

OpenStudy (lukeskylicker):

mk fair enough. So what we have to do is substitute (Sorry for the delay haven't had to do this type of problem in a while had to get my brain remembering this).

OpenStudy (anonymous):

its all good

OpenStudy (lukeskylicker):

We know that the values of x and y will be identical for both equations.

OpenStudy (anonymous):

okay

OpenStudy (lukeskylicker):

x + y =-1 x - 3y =11 What that means is we have to take one these equations and substitute for x or y in the other equation. Or perhaps more simply we need to isolate one variable on one side and substitute the isolated variable for it in the other equation.

OpenStudy (anonymous):

alright

OpenStudy (lukeskylicker):

Let's use the second equation since it isn't just the variables it has that lovely 3y.

OpenStudy (anonymous):

k

OpenStudy (lukeskylicker):

So what we must do is cancel out 3y.

OpenStudy (lukeskylicker):

We do this by using the reverse of the number/variable.

OpenStudy (lukeskylicker):

If I had -3 for example what number would I have to add to get 0.

OpenStudy (lukeskylicker):

You there?

OpenStudy (anonymous):

ya sry you add three

OpenStudy (lukeskylicker):

Exactly.

OpenStudy (lukeskylicker):

Any reverse of the same number always equals 0

OpenStudy (anonymous):

yah

OpenStudy (lukeskylicker):

So -3y + 3y would equal 0. However due to the law of symmetry we must do it to both sides of the equation. So our second equation looks like this: x - 3y + 3y = 11 + 3y however since were canceling. We can safely remove the 3y's on the left side of the equation thus giving us x = 11 + 3y

OpenStudy (lukeskylicker):

Think you understand so far?

OpenStudy (anonymous):

sorta

OpenStudy (lukeskylicker):

OK now here is the tricky part.

OpenStudy (lukeskylicker):

No we must take our second equation. and substitute it for x in the first equation.

OpenStudy (anonymous):

mk

OpenStudy (anonymous):

okay

OpenStudy (lukeskylicker):

Sorry I mean x + y = -1 is now 11 + 3y + y = -1 Typoed.

OpenStudy (anonymous):

its ok

OpenStudy (lukeskylicker):

So now let's add like terms. 3y + y is equal to 4y.

OpenStudy (lukeskylicker):

So our equation is now 11 + 4y = -1

OpenStudy (lukeskylicker):

Now we do the same that did for the second equation. Think you can do this one?

OpenStudy (anonymous):

so im going to add like nterms

OpenStudy (lukeskylicker):

No we've already done that. What were going to do is cancel out terms on the left side.

OpenStudy (lukeskylicker):

However. Keep in mind we do the reverse of PEMDAS

OpenStudy (lukeskylicker):

In other words we do addition and subtraction cancellation instead of multiplication and division cancellation first.

OpenStudy (lukeskylicker):

So we have 11 + 4y = -1 so let's get rid of 11 first.

OpenStudy (lukeskylicker):

subtract 11 from both sides and we get 11 - 11 + 3y = -1 - 11

OpenStudy (lukeskylicker):

or 4y = -12

OpenStudy (anonymous):

my head just exploded

OpenStudy (lukeskylicker):

stay with me. Were almost done.

OpenStudy (anonymous):

okay

OpenStudy (lukeskylicker):

Alright. So with 4y we are multiplying y 4 times. or in other words we have y + y + y + y

OpenStudy (lukeskylicker):

or rather adding y not multiplying.

OpenStudy (lukeskylicker):

So how do we cancel 4y.

OpenStudy (anonymous):

we add 4y to both sides?

OpenStudy (lukeskylicker):

No we do not. What is the reverse of multiplication.

OpenStudy (anonymous):

division

OpenStudy (lukeskylicker):

There you go. So our equation is 4y / 4y = -12 / 4y

OpenStudy (lukeskylicker):

or actually hang on a second. we don't divide by 4y that would just leave us with a 1 on the left side. What we do is is 4y / 4 = -12 / 4

OpenStudy (lukeskylicker):

Or in other words. y = -3

OpenStudy (lukeskylicker):

So now we know the value of y.

OpenStudy (anonymous):

ok

OpenStudy (lukeskylicker):

Now we need to find the value of x. But since we know the value of y it will be a lot easier to find the value of x.

OpenStudy (lukeskylicker):

So we take any of the two equations (let's do the first one and substitute y for -3

OpenStudy (lukeskylicker):

x + y = -1 is now x + -3 = -1

OpenStudy (lukeskylicker):

We do cancellation one last time and we get x - 3 + 3 = -1 + 3

OpenStudy (lukeskylicker):

Or x = 2

OpenStudy (lukeskylicker):

and that is the awnser for your question. x = 2 and y = -3 or (2, -3).

OpenStudy (lukeskylicker):

I recommend bookmarking this page if you need a refresher for this.

OpenStudy (anonymous):

okay thank you

OpenStudy (lukeskylicker):

It's a pleasure.